OpenStudy (anonymous):

Ben started a job that paid $40,000 a year. Each year after the first, his salary was increased by 4%. a. what was Ben's salary in his eight year of employment? b. What is the total amount that Ben earned in eight years?

hero (hero):

First find the salary for each year starting with 1st = 40,0000 2nd = 3rd = . . . 8th = Then use geometric series formula

OpenStudy (anonymous):

ok, give me a sec

OpenStudy (anonymous):

its not arithmetic

OpenStudy (anonymous):

oh you changed it, never mind

OpenStudy (anonymous):

do you know what 1.04^7 is, i dont have a calculator...

OpenStudy (anonymous):

Am i doing it right?

OpenStudy (anonymous):

does anyone know what 1.04^7 is?

OpenStudy (anonymous):

is the answer 5236.2711696?

OpenStudy (anonymous):

DUDEEEEE ANSWER MEEEEEE

OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

@LonelyandForgotten

OpenStudy (ghazi):

have you solved it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my answer is above

OpenStudy (ghazi):

8 years ago , his salary was 28800

OpenStudy (ghazi):

annual

OpenStudy (anonymous):

ago?

OpenStudy (anonymous):

how did u get that

OpenStudy (ghazi):

this is what your question says :) 8 years ago

OpenStudy (anonymous):

where does it say that?

OpenStudy (ghazi):

i guess i misread or misunderstood the question, the first part

OpenStudy (anonymous):

its alright, please try to find the answer and get back to me

OpenStudy (ghazi):

sure :)

hero (hero):

No @ghazi, you mis-interpreted the question. It says nothing about "eight years ago".

OpenStudy (ghazi):

@hero yeah that's what i said :(

OpenStudy (anonymous):

does anyone understand it yet? Still need to understand it if you guys can get back to me

OpenStudy (anonymous):

i gave an answer above, was it right?

OpenStudy (anonymous):

is it correct that i do 1.04^7?

OpenStudy (anonymous):

because thats the common ratio

OpenStudy (anonymous):

and n becomes 7

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

its neither arithmetic, nor geometric, its something else only..... 1st term a1= 40000 2nd term a2= 40000+0.04*40000 = 1.04a1 3rd term a3 = a2+0.04a2 = 1.04 a2 and so on. an = 1.04*a(n-1) with a1=40000

hero (hero):

@hartnn, I hate to disagree with you, but it is geometric

hartnn (hartnn):

oh...but how ?

hero (hero):

I'm not going to try to explain. Look up the general formula for sum of geometric series

OpenStudy (anonymous):

its because theres a common ratio, im pretty sure

hartnn (hartnn):

1) Each year after the first, his salary was increased by 4% 'of first yr salary' <-----------yes , its arithmetic series with common diff. 0.04*40000 2) Each year after the first, his salary was increased by 4%. 'of previous yr' <-----------neither geometric nor arithmetic...and general term is as i shown above.

hartnn (hartnn):

an = 1.04*a(n-1) an = 1.04^(n-1)*a1 so after 8 years a8 = 1.04^7 a1 = 1.04^7 * 40000= .....

hartnn (hartnn):

^ that was for 2nd case. for 1st case , u directly have formulas for arithmetic series.

OpenStudy (anonymous):

the second part of the question asks for the total amount earned over the 8 years which means you use the sum formula, s_n/2(a1+an)? Correct me if im wrong

hartnn (hartnn):

i would like to correct myself : 1) Each year after the first, his salary was increased by 4% 'of first yr salary' <-----------yes , its arithmetic series with common diff. 0.04*40000 2) Each year after the first, his salary was increased by 4%. 'of previous yr' <------------geometric ...and general term is as i shown above. to answer that Q of yours, what should i assume ? 1st case or 2nd case ?

hartnn (hartnn):

i think 2nd case, as correctly pointed by Hero, that its geometric. and you have standard formula for sum of Geometric series.

OpenStudy (anonymous):

I just wanna know if my answer for the first case if correct, i got 5236.2711696

hartnn (hartnn):

how its less than 40 000 ?? 1.04^7 * 40 000 = 52637.27 that will be his salary after 8 years.

hartnn (hartnn):

for sum, use formula, a(r^n -1) / (r-1) r= 1.04, n=8,a=40 000

OpenStudy (anonymous):

my mistake, i got that on my calculator though typed it in wrong

hartnn (hartnn):

no problem! even i did a mistake interpreting this problem at first, seems Hero was correct all along....